# Harmonic Tutte polynomials of matroids III

### v1

@@ 5. Applications to subcode designs
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. Let $r,i$ be the positive integers such that $r\leq k$ and $i\geq d_{r}$. Then the set $\mathcal{S}_{r,i}(C)$ forms a $t$-design if and only if $W_{C,f}^{(r)}(x,y)=0$ for all $f\in\operatorname{Harm}_{d}(n)$, $1\leq d\leq t$.
@@
$\mathcal{L}_{r,t}:=\{i\in\{d_{r}^{\perp},\ldots,n-t\}\mid A_{i}^{(r)}(C^{\perp})\neq 0\},$
@@
Let $f\in\operatorname{Harm}_{d}(n)$ for $1\leq d\leq t$. By Lemma 5.1, it is sufficient to show that $A_{i,f}^{(r)}(C)=0$ (resp. $A_{i,f}^{(r)}(C^{\perp})=0$) for all $i=d,\ldots,n-d$. Applying Theorem 3.1 to the dual code $C^{\perp}$ gives the identity
(5) $\sum_{\ell=1}^{r}(-1)^{d}M_{\ell}^{r}\underline{A}^{(\ell)}(C^{\perp};f)=\underline{A}^{(r)}(C;f),$
@@
where $k^{\perp}$ denotes the dimension of $C^{\perp}$ and $K_{p}(i-d;n-2d,q^{j})$ is the Krawtchouk polynomial defined by (see [23])
Since $d_{r}$ is the $r$-th generalized Hamming weight of $C$,
@@ Acknowledgements
The second named author is supported by SUST Research Centre under Research Grant PS/2023/1/22. The third named author is supported by JSPS KAKENHI (22K03277).

### v2

@@ 5. Applications to subcode designs
Example 5.1.
Let $C$ be a $[6,3]$ code over $\mathbb{F}_{2}$ with generator matrix
$\begin{pmatrix}1&1&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&1&1\end{pmatrix}.$
There are seven linear subcodes of $C$ with dimension $2$. Three of them have weight $4$; the set of supports of these three is
$\mathcal{S}_{2,4}=\{\{1,2,3,4\},\{1,2,5,6\},\{3,4,5,6\}\}\,.$
Note that $\mathcal{S}_{2,4}(C)$ is a $1$-design. Moreover, for $f\in\operatorname{Harm}_{1}(6)$, where $f=a\{1\}+b\{2\}+c\{3\}+d\{4\}+e\{5\}-(a+b+c+d+e)\{6\}$, we get $A_{4,f}^{(2)}(C)=0$.
@@
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. Let $r,i$ be the positive integers such that $r\leq k$ and $i\geq d_{r}$. Then the set $\mathcal{S}_{r,i}(C)$ forms a $t$-design if and only if $A_{i,f}^{(r)}(C)=0$ for all $f\in\operatorname{Harm}_{d}(n)$, $1\leq d\leq t$.
@@
$\mathcal{L}_{r,t}:=\bigg{\{}i\in\{d_{r}^{\perp},\ldots,n-t\}\;\bigg{|}\;{\sum_{\ell=0}^{r}[r]_{\ell}\,A_{i}^{(\ell)}(C^{\perp})}\neq 0\bigg{\}},$
@@
Let $f\in\operatorname{Harm}_{d}(n)$ for $1\leq d\leq t$. By Lemma 5.1, it is sufficient to show that $A_{i,f}^{(r)}(C)=0$ (resp. $A_{i,f}^{(r)}(C^{\perp})=0$) for all $i=d,\ldots,n-d$. Applying Theorem 3.1 and Remark 4.2 to the dual code $C^{\perp}$ gives the identity
(6) $(-1)^{d}\sum_{\ell=1}^{r}M_{\ell}^{r}\underline{A}^{(\ell)}(C^{\perp};f)=\underline{A}^{(r)}(C;f),$
@@
where $k^{\perp}$ denotes the dimension of $C^{\perp}$ and $K_{p}(i-d;n-2d,q^{j})$ is the Krawtchouk polynomial given in Remark 4.1. Since $d_{r}$ is the $r$-th generalized Hamming weight of $C$,
@@ Acknowledgements
This work was supported by JSPS KAKENHI (22K03277) and SUST Research Centre (PS/2023/1/22). The authors would also like to thank the anonymous reviewers for their beneficial comments on an earlier version of the manuscript.